Circular No. 6663

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970508 I. A. Smith, Rice University; and R. A. Gruendl, University of Illinois

at Urbana/Champaign report, on behalf of the Berkeley-Illinois-Maryland Association (BIMA) GRB team: “Using the BIMA array at 3.5 mm, we mapped most of the 5′ γ-ray-burst error box reported on IAUC 6649 for 5.5 hr, starting May 10.1 UT. This region includes the new x-ray source SAX J0653.8+7916 (IAUC 6656). No obvious sources were detected, with a 1σ rms of 10 mJy. A deeper stare of SAX J0653.8+7916 was performed between May 10.8 and 11.1, although 2 hr were lost because of storms. No sources were detected, with a 1σ rms of 2.0 mJy. In another stare between May 15.1 and 15.3, there may be a mm-continuum source at the location of the radio variable reported on IAUC 6662, though our preliminary analysis gives a flux of 5.5 ± 1.7 mJy, which is only marginally significant.”

C. Chevalier and S. A. Ilovaisky, Observatoire de Haute-Provence, report that three 15-min CCD exposures of Bond’s variable object (IAUC 6654), obtained with the 1.2-m Haute-Provence telescope (+ Cousins R filter) on May 14.40 UT under nonphotometric conditions, shows that it has further decreased in brightness, to R = 21.1 ± 0.2 (using as comparison the magnitude given by Schaefer et al. on IAUC 6658 for the brighter star located 13′′ to the north).

A. I. Kopylov, V. V. Sokolov, and S. V. Zharikov, Special Astrophysical Observatory, Russian Academy of Sciences; and M. Feroci, E. Palazzi, and the rest of the BeppoSAX GRBM team, report: “Optical multicolor observations of Bond’s variable object (IAUC 6654) were obtained with 1-m (May 9.89 UT) and 6-m telescopes. Our preliminary reduction shows the following Johnson and Kron-Cousins magnitudes: May 9.89, R_{c} = 21.1; May 10.77, B = 20.90, V = 20.34, R_{c} = 19.87, I_{c} = 19.43; 10.93, 20.79, 20.33, 19.92, 19.30; 11.76, 21.42, 20.77, 20.18, 19.48; 12.87, 21.70, 21.19, 20.77, 20.25; 13.88, 21.93, 21.33, 21.01, 21.01. Estimated errors are ∼ 0.15 for the 1-m telescope measurements and 0.05–0.08 for the others. These measurements of the ‘broadband spectrum’ roughly correspond to a power law with a spectral index of −1.5 for May 10 and 12, and −2.2 for May 11. During a 1-hr interval on May 13, the object showed variability of ∼ 0.2 mag both in the B and V bands, though only one exposure was obtained in the I_{c} band. The unusual ‘Z-shaped’ object (IAUC 6658) is resolved into two components in the May 13 R_{c} image (FWHM seeing = 1′′.6): a very red point source (R_{c} = 21.7) adjacent southward to an arc-like extended source of size 5′′.5 from tip to tip and 2′′–2′′.5 in width. The brightness of the extended source increases from the middle to the tips.”

1997 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green